OpenStudy (anonymous):

I am trying to do Direct Linear Variations and i am not sure how. can someone help me with like 5 questions? 1: The number of calories burned, c, varies directly with the time spent exercising, t. When Lila bikes for 3 hours, she burns 900 calories. Which of the following equations shows this direct linear variation? A. c = 300t B. c = 3t C. c = 900t D. c = t i think A or C 2: Amelia paid $252 for 28 cases of granola bars for the snack bar. The total cost varies directly with the number of cases purchased. How much will she pay for 10 cases of granola bars

OpenStudy (prettygirl_shynice):

Welcome to open study how are you

OpenStudy (prettygirl_shynice):

okay you ready

OpenStudy (anonymous):

3: A customer paid $4.50 for 6 photo prints. The cost for prints varies directly with the number of prints made. How much will the customer pay for 15 prints? 4: A horse runs for 0.25 h and travels 9 mi. The horse’s distance varies directly with the amount of time it runs. How many miles will it run in 1.5 h? 5: A wheel rotates on an axle 84 times in 6 min. The number of rotations varies directly with the time the wheel rotates. How long will it take the wheel to rotate 168 times?

OpenStudy (anonymous):

im ready lol

OpenStudy (prettygirl_shynice):

thats Let y be the number of rotations Let x be the time in minutes 84 = 6k y = k x ( y varies directly with x) 84 = 6k k = 14 y = 14 x 168 = 14 x when y=168 , x = ? x = 12 minutes number 5 thats

OpenStudy (prettygirl_shynice):

@lillieJULE

OpenStudy (anonymous):

okay that makes alot more sense than what they tried teaching me

OpenStudy (prettygirl_shynice):

OK

OpenStudy (prettygirl_shynice):

REDY FOR NUMBER 4 AND 3

OpenStudy (prettygirl_shynice):

0.5 hrs=18 miles so 1 hr=36 miles, 36+18=? that 4

OpenStudy (prettygirl_shynice):

@lillieJULE got it

OpenStudy (anonymous):

yep

OpenStudy (prettygirl_shynice):

Well you would divide 4.50 ÷ 6 so .75 then .75 x 15. So 11.25 thats number 3

OpenStudy (prettygirl_shynice):

thats the last one

OpenStudy (prettygirl_shynice):

@lillieJULE that last one

OpenStudy (anonymous):

wat about 2 and 1?

OpenStudy (anonymous):

@Prettygirl_shynice

OpenStudy (prettygirl_shynice):

yea

OpenStudy (anonymous):

im don't see 2 and 1 only 5, 4, 3

OpenStudy (prettygirl_shynice):

okay im doiung em

OpenStudy (anonymous):

kk thanks

OpenStudy (prettygirl_shynice):

5.the answer is B. C= 300t. I just had a math question like this yesterday. :) C= 300 * 3 = 900.

OpenStudy (prettygirl_shynice):

1.the answer is B. C= 300t. I just had a math question like this yesterday. :) C= 300 * 3 = 900.

OpenStudy (anonymous):

lol k

OpenStudy (prettygirl_shynice):

2.This question requires you find the total cost per case 28x=252 divide 28 from each side x=9 Multiply 9 by 10 to get your answer of $90

OpenStudy (prettygirl_shynice):

ok

OpenStudy (anonymous):

thnx

OpenStudy (prettygirl_shynice):

your welocme tell me what your scrose is

OpenStudy (prettygirl_shynice):

ment tell me what your score is

OpenStudy (prettygirl_shynice):

@lillieJULE